# Framing in the Presence of Supporting Data: A Case Study in U.S. Economic News

### v2

@@ 5.1 Frame Prediction
We perform 5-fold cross-validation in all scenarios, and assume a multi-class setup for each frame component. To create our folds, we first split the articles that were cross-annotated (270 articles, 689 quantities) into five folds, and consider 4 folds for training and 1 for testing. Then, we further enhance the training data for all cases with the additional annotated articles, which we consider as a source of noisy supervision. More details about the data splits used for experiments are included in App. A.6.
@@ Results
We present our general results in Tab.  3. We find that DAPT is considerably helpful for quantity-level predictions, but does not improve article-level predictions. We hypothesize that it is difficult for RoBERTa to model the long context of articles to take advantage of the masked language modeling objective. Next, we find that the relational model gives us a clear boost in performance over the base classifiers, with the exception of article type. This improvement supports our modeling decision to decompose frame prediction into inter-dependent sub-tasks. While Tab. 3 reports macro F1, we find the same trend for weighted F1. These results are included in App. A.7.
@@
We include results for the best relational model by article publisher in Tab. 5. At the article level, we observe higher values for the New York Times, reflecting the higher presence of annotated examples for this publisher. Interestingly, the results for Breitbart and Huffpost do not follow this trend, maintaining relatively high values in spite of less supervision. In the case quantities, we see relatively high and stable performance for all publishers, with the exception of Fox News. This is particularly encouraging, as it suggests that we can leverage quantity predictions to perform a high-fidelity selection and framing analysis of economic news.
@@
Lastly, we include fine-grained results for predicting the different macro-indicators using our best model in Tab. 6. We find that we have generally good performance for most, with the exception of the ’other’ category, which may be due to the lack of cohesiveness in this class. We also see lower performance for retail sales, which we attribute to the lack of sufficient supervision for this class (See Fig. 2). An effective large-scale analysis of retail indicators in the news would require greater supervision for this class.
@@ 5.2 Framing and Selection Analysis
Our preliminary evidence suggests that the framing of articles on any given topic by any given publication is both sensitive to exogenous factors as well as decisions made by the publication itself.
Figure 3: Framing of articles referencing job numbers from 2015 through 2023 in the New York TimesThe Y-axis has been limited to account for the massive swing in employment during the early pandemic.
Figure 4: Selection of economic indicators referencing price (price & energy) numbers from 2015 through 2023 in the New York Times, Washington Post, and Wall Street Journal
Fig. 3 illustrates how the New York Times (NYT) framed their articles in regard to jobs data from 2015 to 2023. First, we can see that prior to 2020, during a period of stable job growth, the NYT already had a sustained negative valance, matching the key findings in recent literature on the general negativity bias on the Mainstream Media van Binsbergen et al. (2024). In response to the onset of COVID in 2020, the NYT intensified its coverage of the job market in response to a dramatic surge in unemployment caused by a cascade of lockdowns. We can also see that much of this coverage carried an extreme negative valence, which matches the large drop in jobs and subsequent uncertainty. But, as the job market recovered and the labor market entered an extended period of exceptionally high growth, the volume of reporting in the NYT decreased back to its pre-pandemic levels, but the negativity persisted. This was not the same for other publications such as the Wall Street Journal, whose coverage more closely resembles the shifts in valence in the supporting data (See App. A.8.)
@@
Fig. 4 shows us another lever that publications can pull when shaping their coverage: indicator selection. As inflation began to rise, we can see that all three publications responded by increasing their coverage of prices through a greater use of price based indicators.This includes energy (e.g. oil) prices. These three publications did not, however, respond in an identical manner, with, for example, the New York Times increasing their coverage of prices far after the other publications returned to pre-crisis levels of coverage.
Between these two figures we can see how through even simple analyses, we can begin to to show that editors have considerable leeway when choosing what indicator to cover at any given time, and how to cover it. While we kept the analysis brief due to space considerations, we were able to show the usability of our framework to track a diverse set of indicators reported in different publishers across time.
@@ 6 Conclusion and Future Work
In this paper, we presented a novel framework for modeling framing in news articles in the presence of supporting data. We focused our analysis on news about the U.S. economy, and proposed an annotation schema to identify framing at three levels of abstraction: the general framing of the economic conditions, the economic indicators that are highlighted, and the framing of each of those indicators. We showed that we can predict the components of our schema with relatively good performance, even in very low supervision settings. Finally, we demonstrated how our framework can be used to perform a large scale exploration of the frame choices for a given economic indicator. We are able to show that different publications cover economics using different indicators, for example the NYT covers jobs numbers less than normal during a strong jobs recovery, instead switching to more coverage rising prices. And, while all three publications responded to rising inflation, they each used price numbers differently when covering the economy, with the NYT employing price numbers more frequently than the Washington Post, or the Wall Street Journal.
Going forward, we have two main points of focus. First, we are working on expanding our schema to capture a larger set of economic aspects. While we focused our analysis on macro-economic indicators, there are other indicators that we could track to get a more holistic view of the economy. For example, for governmental indicators, we could track types of expenditures (e.g. social security), revenue (e.g. taxes), as well as debt and deficit. Additionally, we want to perform a large-scale analysis of economic news framing. In this paper, we presented a brief demonstration of how our framework can be used to analyze shifts in framing across time for different publications. Next, we want to perform a similar analysis for a larger set of economic indicators and publications. We note that to perform such an analysis we must scale up our annotations to improve our model predictions.
@@ 7 Limitations
The work presented in this paper has two main limitations. (1) Obtaining high-quality annotations for our frame structure is very expensive and time consuming. For this reason, we worked with a low amount of supervision. While we showed that we could obtain relatively good performance in this constrained scenario, to be able to use our framework for a large scale, holistic analysis, we need to address this issue. This is particularly important given that our annotations are considerably skewed. For example, we have significantly more supervision for job numbers and market numbers (which occur more often) than we do for energy prices (which occur less often). This considerably affects our performance for more long-tail indicators, and hence limits the types of analysis that we can perform. Our current efforts are dedicated to increase the amount of annotations for all components of our frame structure, as well as incorporating alternative semi-supervised strategies. We are hoping to release an extended version of the annotated dataset to the community in the next few months. (2) The fact that we are using automated techniques for frame prediction necessarily carries some uncertainty. Even if we improve our models considerably, our large scale analysis has a margin of error. It is important to acknowledge this when presenting our findings.
@@ 8 Ethical Considerations
To the best of our knowledge, no code of ethics was violated during the development of this project. We used publicly available tools to collect our dataset, and discarded any instances that were unreachable. The annotators were paid \$15 per hour, and no personally identifiable information was collected or recorded during annotation. We did a thorough evaluation of our dataset, which is presented in the paper. We reported all pre-processing steps, learning configuration, hyperparameters, and additional technical details. In the interest of reproducibility, the data and code will be released to the community. Due to space constraints, some information was relegated to the Appendix. The results reported in this paper support our claims and we believe that they are reproducible. The analysis reported in Section 5.2 is done using the outputs of a machine learning model and do not represent the authors personal views. The uncertainty of these predictions was adequately acknowledged in the Limitations Section, and the estimated accuracy was reported in Section 5.1.
@@ A.1 Additional Data Collection Details
We stored the document’s title as extracted by readability and regarded it as the article’s headline. In some cases, there were errors when extracting headlines. To deal with these cases, we used T5 Raffel et al. (2020), a generative language model, to generate headlines using the body of the article. To do this, we used a model that was trained on 500,000 articles with their headlines ******huggingface.co/Michau/t5-base-en-generate-headline.
To identify the subset of articles that discuss the economy, we curated a lexicon of economic terms, shown in Table 7. We sourced these terms from the Bureau of Labor Statistics and the FRED database operated by the Federal Reserve bank of St. Louis.††††††https://www.bls.gov/data/‡‡‡‡‡‡https://fred.stlouisfed.org/ These sources provide a fairly comprehensive list of common metrics of economic activity.
@@ A.2 Codebook Development
We inherited the type category from the indicators to codify the “prevalent” or “dominant” type of information that the article discussed. For the economic conditions and direction, we took the Gallup economic index questions******https://news.gallup.com/poll/1609/consumer-views-economy.aspx, and re-worded them to reflect the fact that coders were reading an article from sometime in the last 8 years. Finally, we refined them after a few rounds of annotation, taking annotator feedback into account.
@@ A.6 Additional Details for Experimental Setup
Classifiers architecture details
@@ A.7 Additional Results for Frame Prediction
A.8 Additional Framing Analysis Graphs

### v3

@@ 5.1 Frame Prediction
To determine the effectiveness of exploiting inter-dependencies with statistical relational learning and our additional pre-training steps, we conducted an ablation study in addition to evaluating the model as a whole. We provide a discussion of the results of these experiments, in which we detail per-class and per-publisher metrics.
@@
We perform 5-fold cross-validation in all scenarios, and assume a multi-class setup for each frame component. To create our folds, we first split the articles that were cross-annotated and where inter-annotator agreement was reached (270 articles, 689 quantities) into five folds, and consider 4 folds for training and 1 for testing. Then, we further enhance the training data for all cases with the additional annotated articles, which we consider as a source of noisy supervision. More details about the data splits used for experiments are included in App. A.6.
@@ Results
We present our general results in Tab.  3. This includes baselines obtained by zero-shot and two-shot prompting Mistral-7b-Instruct-v0.2. Additional details about the prompts and performance can be found in App. A.6. We find that fine-tuning, even with relatively little data, results in better performance than prompting an LLM. Additionally, we find that DAPT is considerably helpful for quantity-level predictions, but does not improve article-level predictions. We hypothesize that it is difficult for RoBERTa to model the long context of articles to take advantage of the masked language modeling objective. Next, we find that the relational model gives us a clear boost in performance over the base classifiers, with the exception of article type. This improvement supports our modeling decision to decompose frame prediction into inter-dependent sub-tasks. While Tab. 3 reports macro F1, we find the same trend for weighted F1 (See App. A.7).
@@
We include results for the best relational model by article publisher in Tab. 5. At the article level, we observe higher values for the New York Times, reflecting the higher presence of annotated examples for this publisher (see Tab.  1). The results for Fox News articles are lower, reflecting the lower number of examples. Interestingly, the results for Breitbart and Huffpost do not follow this trend, maintaining relatively high values in spite of less supervision. In the case of quantities, we see relatively high and stable performance for all publishers, with the exception of Fox News. We provide further error analysis in regard to this in App.  A.9. These results are particularly encouraging, as they suggest that we can leverage quantity predictions to perform a high-fidelity selection and framing analysis of economic news.
@@
Lastly, we include fine-grained results for predicting the different macro-indicators using our best model in Tab. 6. We find that we have relatively good performance for most indicators, with the exception of the ’other’ category, which may be due to the lack of cohesiveness in this class. We also see lower performance for retail sales, which we attribute to the lack of sufficient supervision for this class (See Fig. 3). An effective large-scale analysis of retail indicators in the news would require greater supervision for this class.
@@ 5.2 Framing and Selection Analysis
Our preliminary evidence suggests that the framing of articles on any given topic by any given publication is both sensitive to exogenous factors as well as decisions made by the publication itself. Fig. 4 illustrates how the New York Times (NYT) framed their articles in regard to jobs data from 2015 to 2023. First, we can see that prior to 2020, during a period of stable job growth, the NYT already had a sustained negative valance, matching the key findings in recent literature on the general negativity bias on the Mainstream Media van Binsbergen et al. (2024). In response to the onset of COVID in 2020, the NYT intensified its coverage of the job market in response to a dramatic surge in unemployment caused by a cascade of lockdowns. We can also see that much of this coverage carried an extreme negative valence, which matches the large drop in jobs and subsequent uncertainty.
As the job market recovered and the labor market entered an extended period of exceptionally high growth, the volume of reporting in the NYT decreased back to its pre-pandemic levels, but the negativity persisted. This was not the same for other publications such as the Wall Street Journal, whose coverage more closely resembles the shifts in valence in the supporting data (See App. A.10.)
Figure 4: Framing of articles referencing job numbers from 2015 through 2023 in the New York Times.The Y-axis has been limited to account for the massive swing in employment during the early pandemic. Spin was aggregated quarterly. Monthly payroll (employment) data can be seen in the dotted black line. The gray bars represent proportion of the overall coverage taken up by jobs reporting.
Figure 5: Selection of economic indicators referencing price (price & energy) numbers from 2015 through 2023 in the New York Times, Washington Post, and Wall Street Journal. Aggregated Quarterly. Monthly CPI data can be seen in the dotted blue line.
@@
Fig. 5 shows us another lever that publications can pull when shaping their coverage: indicator selection. As inflation began to rise, we can see that all three publications responded by increasing their coverage of prices through a greater use of price based indicators.******This includes energy (e.g. oil) prices. These three publications did not, however, respond in an identical manner, with, for example, the New York Times increasing their coverage of prices far after the other publications returned to pre-crisis levels of coverage.
With the simple analysis shown in these two figures, we can begin to show that editors have considerable leeway when choosing what indicator to cover at any given time, and how to cover it. While we kept the analysis brief due to space considerations, we were able to show the usability of our framework to track a diverse set of indicators reported in different publishers across time.
6 Applicability to Other Domains
We emphasize that the main contribution of this paper is not centered around performance gains of our relational model over the baseline models, but rather the operationalization of a challenging task: analyzing framing at a scale for domains with supporting numerical data, such as the economy. While our solution does not apply to all framing scenarios, it applies to a substantial number. For example, we envision this framework being useful for analyzing news about crime (by tracking data points maintained by the FBI and local jurisdictions), climate change (by tracking weather statistics, global avg. temperature, emissions statistics, etc.) and public-opinion (by tracking polling numbers, political betting market numbers, etc.).
@@ 7 Conclusion and Future Work
In this paper, we presented a novel framework for modeling framing in news articles in the presence of supporting data. We focused our analysis on news about the U.S. economy, and proposed an annotation schema to identify framing at three levels of abstraction: the general framing of the economic conditions, the economic indicators that are highlighted, and the framing of each of those indicators.
We showed that we can predict the components of our schema with relatively good performance, even in very low supervision settings. Finally, we demonstrated how our framework can be used to perform a large scale exploration of the frame choices for a given economic indicator.
In our analysis, we showed that different publications cover economics using different indicators, for example the NYT covers jobs numbers less than normal during a strong jobs recovery, instead switching to more coverage rising prices. And, while all three publications responded to rising inflation, they each used price numbers differently when covering the economy, with the NYT employing price numbers more frequently than the Washington Post, or the Wall Street Journal.
Going forward, we have two main points of focus. First, we are working on expanding our schema to capture a larger set of economic aspects. While we focused our analysis on macro-economic indicators, there are other indicators that we could track to get a more holistic view of the economy. For example, for governmental indicators, we could track types of expenditures (e.g. social security), revenue (e.g. taxes), as well as debt and deficit. Additionally, we want to perform a large-scale analysis of economic news framing. In this paper, we presented a brief demonstration of how our framework can be used to analyze shifts in framing across time for different publications. Next, we want to perform a similar analysis for a larger set of economic indicators and publications. We note that to perform such an analysis we must scale up our annotations and improve our model predictions.
@@ 8 Limitations
The work presented in this paper has two main limitations:
(1) Obtaining high-quality annotations for our frame structure is very expensive and time consuming. For this reason, we worked with a low amount of supervision. While we showed that we could obtain relatively good performance in this constrained scenario, to be able to use our framework for a large scale, holistic analysis, we need to address this issue. This is particularly important given that our annotations are considerably skewed. For example, we have significantly more supervision for job numbers and market numbers (which occur more often) than we do for energy prices (which occur less often). This considerably affects our performance for more long-tail indicators, and hence limits the types of analysis that we can perform. Our current efforts are dedicated to increase the amount of annotations for all components of our frame structure, as well as incorporating alternative semi-supervised strategies. We are hoping to release an extended version of the annotated dataset to the community in the near future.
(2) The fact that we are using automated techniques for frame prediction necessarily carries some uncertainty. Even if we improve our models considerably, our large scale analysis will have a margin of error. It is important to acknowledge this when presenting our findings.
@@ 9 Ethical Considerations
To the best of our knowledge, no code of ethics was violated during the development of this project. We used publicly available tools to collect our dataset, and discarded any instances that were unreachable. The annotators were paid \$15 per hour, and no personally identifiable information was collected or recorded during annotation.
We performed a thorough evaluation of our dataset, which is presented in the paper. We reported all pre-processing steps, learning configurations, hyperparameters, and additional technical details. Due to space constraints, some of this information was relegated to the Appendix. Further, the data and code have been released to the community. The results reported in this paper support our claims and we believe that they are reproducible.
The analysis reported in Section 5.2 is done using the outputs of a machine learning model and does not represent the authors personal views. The uncertainty of these predictions was adequately acknowledged in the Limitations Section, and the estimated accuracy was reported in Section 5.1.
10 Acknowledgements
This work utilized the Alpine high-performance computing resource, the Blanca condo computing resource, and the CUmulus on-premise cloud service at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, Colorado State University, and the NSF (award 2201538). Blanca is jointly funded by computing users and the University of Colorado Boulder. CUmulus is jointly funded by the National Science Foundation (award OAC-1925766) and the University of Colorado Boulder.
@@ A.1 Additional Data Collection Details
We stored the document’s title as extracted by readability and regarded it as the article’s headline. In some cases, there were errors when extracting headlines. To deal with these cases, we used T5 Raffel et al. (2020), a generative language model, to generate headlines using the body of the article. To do this, we used a model that was trained on 500,000 articles with their headlines ††††††huggingface.co/Michau/t5-base-en-generate-headline.
To identify the subset of articles that discuss the economy, we curated a lexicon of economic terms, shown in Table 7. We sourced these terms from the Bureau of Labor Statistics and the FRED database operated by the Federal Reserve bank of St. Louis.‡‡‡‡‡‡https://www.bls.gov/data/******https://fred.stlouisfed.org/ These sources provide a fairly comprehensive list of common metrics of economic activity.
@@ A.2 Codebook Development
We inherited the type category from the indicators to codify the “prevalent” or “dominant” type of information that the article discussed. For the economic conditions and direction, we took the Gallup economic index questions*†*†*†https://news.gallup.com/poll/1609/consumer-views-economy.aspx, and re-worded them to reflect the fact that coders were reading an article from sometime in the last 8 years. Finally, we refined them after a few rounds of annotation, taking annotator feedback into account.
@@ A.6 Additional Details for Experimental Setup
Mistral baseline details
To further motivate our modeling choices, we present baseline results for zero and few-shot prompting instruction tuned model, Mistral-7B-Instruct-v0.2, to obtain labels for our dataset. Below we include samples of the zero-shot prompts used to obtain these results.
Example prompt for article-level type prediction
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy:
So for instance the following:
excerpt: [article text]
multiple choice question: What is the main type of economic information covered in this article?
A. Firm-specific
B. Industry-specific
C. Macroeconomic / General Economic Conditions
D. Government revenue and expenses
E. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
Example prompt for quantity-level type prediction
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy.
So for instance the following:
excerpt: [indicator text]
context: [context text]
multiple choice question: The excerpt should contain an economic indicator value. Based on the context, what type of indicator is it?
A. Macroeconomic / General Economic Conditions
B. Industry-specific
C. Government revenue and expenses
D. Personal
E. Firm-specific
F. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
The prompts were used for each sample in the test set of each fold. For two-shot prompting, two examples were randomly selected from the training set and enveloped into the prompt as shown below.
Example two-shot prompt for article-level type prediction
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy:
So for instance the following:
excerpt: [example 1 article text]
multiple choice question: What is the main type of economic information covered in this article?
A. Firm-specific
B. Industry-specific
C. Macroeconomic / General Economic Conditions
D. Government revenue and expenses
E. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
{‘role’: ‘assistant’, ‘content’: [example 1 gold label]’ }
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy:
So for instance the following:
excerpt: [example 2 article text]
multiple choice question: What is the main type of economic information covered in this article?
A. Firm-specific
B. Industry-specific
C. Macroeconomic / General Economic Conditions
D. Government revenue and expenses
E. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
{‘role’: ‘assistant’, ‘content’: [example 2 gold label]’ }
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy:
excerpt: [article text]
multiple choice question: What is the main type of economic information covered in this article?
A. Firm-specific
B. Industry-specific
C. Macroeconomic / General Economic Conditions
D. Government revenue and expenses
E. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
Example two-shot prompt for quantity-level type prediction
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy.
So for instance the following:
excerpt: [example 1 indicator text]
context: [example 1 context text]
multiple choice question: The excerpt should contain an economic indicator value. Based on the context, what type of indicator is it?
A. Macroeconomic / General Economic Conditions
B. Industry-specific
C. Government revenue and expenses
D. Personal
E. Firm-specific
F. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
{‘role’: ‘assistant’, ‘content’: [example 1 gold label]’ }
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy.
So for instance the following:
excerpt: [example 2 indicator text]
context: [example 2 context text]
multiple choice question: The excerpt should contain an economic indicator value. Based on the context, what type of indicator is it?
A. Macroeconomic / General Economic Conditions
B. Industry-specific
C. Government revenue and expenses
D. Personal
E. Firm-specific
F. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
{‘role’: ‘assistant’, ‘content’: [example 2 gold label]’ }
{‘role’: ‘user’, ‘content’: ‘You are a helpful annotation assistant. Your task is to answer a multiple choice question based on the below information from a U.S. news article about the economy.
excerpt: [indicator text]
context: [context text]
multiple choice question: The excerpt should contain an economic indicator value. Based on the context, what type of indicator is it?
A. Macroeconomic / General Economic Conditions
B. Industry-specific
C. Government revenue and expenses
D. Personal
E. Firm-specific
F. None of the above
Please answer with a single letter without explanations. If you are unsure, please guess.’}
Article-level RoBERTa base classifier architecture details
@@ A.7 Additional Results for Frame Prediction
A.8 Addressing Concerns with Data Leakage and DAPT
We chose to include the annotated subset of data for DAPT because, while one may not have access to a labeled dataset, anyone utilizing the framework would have access to their full unlabelled dataset, and thus would be able to perform additional rounds of DAPT using self-supervised objectives.
However, to address any concerns with data leakage, we executed DAPT with all labelled articles removed from the training set. Specifically, any articles used at any point for fine-tuning the classifiers were omitted. We present results for all versions of the model which included DAPT with this checkpoint in Tab. 12. We find that there is no significant advantage to including labelled articles in the pre-training set.
Model Article-level Quantity-level
Type Cond Dir Type Ind +/+/-+ / -
RoBERTa + DAPT 0.382 0.636 0.47 0.697 0.769 0.776
RoBERTa + DAPT (LAR) 0.519 0.521 0.426 0.700 0.722 0.763
RoBERTa + DAPT + all noisy data 0.474 0.574 0.475 0.722 0.817 0.803
RoBERTa + DAPT (LAR) + all noisy data 0.465 0.562 0.429 0.698 0.787 0.776
RoBERTa + DAPT (LAR) + best noisy data point 0.446 0.487 0.382 0.688 0.870 0.821
Table 12: Base models trained with DAPT, labelled articles removed (LAR) from pre-training set
A.9 Error Analysis for Frame Prediction
In Tab. 5, we observe particular instability for predictions over articles from Fox News and the Wall Street Journal.
For Fox News articles, performance for article-level predictions are particularly low. We attribute this to the fact that there are only three training examples in the data. This prohibits the model from learning to make predictions for the slightly different writing style. Additionally, this means that the test set is very small and one incorrect prediction decreases the macro f1 score significantly. We note that the training example for which an incorrect article type was predicted, had a gold label of government but was predicted macro. This was also a common disagreement among annotators, as described in App. A.5, and therefore a more difficult problem.
Quantity-level polarity was also unstable for Fox News. There were only two training examples for polarity. For the single error, the model predicted neutral while the gold label was negative. This was also a common disagreement among annotators.
In spite of sufficient training examples, we achieve relatively low f1 scores on article-level predictions for documents from the Wall Street Journal. We believe that this may be attributed to the difference in format of these articles compared to those of other publishers. In Fig. 14, we observe that of the 18 training examples, 14 had macro gold labels. We believe that the class imbalance caused the model to associate the Wall Street Journal writing with a macro label only. This is supported by the fact that the model predicted that 17 of the 18 articles were of type macro.
Because the relational model predicts None for the other article-level annotation components when a label other than Macro is predicted for article type, this type of error contributes to the lower F1 scores for article-level condition and direction. We believe these errors can also be attributed to the somewhat dryer delivery of information in WSJ articles. Empirically, their articles tend to express less opinion when discussing the economy compared to other publishers in the corpus.
Figure 14: A significant majority of Wall Street Journal training articles were of type macro, causing the model to be biased toward this label.
A.10 Additional Framing Analysis Graphs